OpenStudy (anonymous):

What is the slope of the line that passes through the points (1, 1) and (–1, –5)?

OpenStudy (anonymous):

Try using the slope formula : slope = (y2 - y1) / (x2 - x1) (1,1)...x1 = 1 and y1 = 1 (-1,-5).. x2 = -1 and y2 = -5 now sub slope = (-5-1) / (-1-1) slope = -6/-2 slope = 3 any questions ?

OpenStudy (anonymous):

@kelliegirl33 did u get my message

OpenStudy (anonymous):

oops...I see it now

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

do you need help now ?

OpenStudy (anonymous):

just tag me when you do :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

do you have any questions ms1231988 ?

OpenStudy (anonymous):

No, i got it thanks

OpenStudy (anonymous):

anytime :)